# Delving Deeper Into Astromorphic Transformers

### v2

@@ III-D1 Hebbian Plasticity
Hebbian plasticity is characterized by the sequential adjustment of synaptic weights in accordance with the interconnection pattern between two neurons (pre and postsynaptic neurons) at the synapse. Previous studies [18] have exclusively focused on the plasticity occurring between two neurons neglecting non-linearities and the plasticity that may occur between astrocytes and postsynaptic neurons, which is also involved in the learning process.
Contribution I: In the context of Hebbian plasticity, we underscore the importance of addition of astrocyte-postsynaptic neuron connected Hebbian-inspired learning along with inherent non-linearities. To elaborate, the synaptic weights ($H_{neuron}$) that establish connections between the presynaptic and postsynaptic neurons are acquired through Hebbian plasticity during the write mode. This learning occurs due to the connection between pre and postsynaptic neuron which can be mapped to the synaptic current ($I_{syn}$) from the neuroscience perspective. As discussed previously, $I_{syn}$ is modulated by DSE so, $H_{neuron}$ is inspired from the direct feedback in the tripartite synapse. This leads to Hebbian plasticity between the hidden layer (presynaptic neuron) and the output layer (postsynaptic neuron). Hidden neuron outputs are denoted as $h_{t}=\phi(k_{t})$ whereas the output (last) layer outputs are denoted as $l_{t}=v_{t}$ in the write mode (refer to Equation 8). $H_{neuron}$ weights are learnt via Equation 9, where $m$ acts as a scaling factor affecting the learning speed.
@@
The astrocyte functions as the third constituent in the tripartite synapse, enabling bidirectional communication with the synaptic neurons. Hence, we propose that the connection between the astrocyte and the postsynaptic neuron be also represented in this formulation through a learnable weight matrix ($H_{astro}$), in accordance with the Hebbian rule. Equation 10 shows the learning mechanism of these weights in which the presynaptic neuron output ($h_{t}$) is replaced by the astrocytic activity, represented by $W_{astro}$, as this connection incorporates the astrocyte and the postsynaptic neuron. The formulation of $W_{astro}$ will be considered in the succeeding text.
@@
The weight, denoted as $H_{astro}$, is inspired from the extrasynaptic current $SIC$ (refer to tripartite synapse dynamics section). According to Equation 7, $SIC$ is added to $I_{syn}$ to generate the total synaptic current. Consequently, the two distinct learnable weights, denoted as $H_{neuron}$ and $H_{astro}$, are combined to produce the Hebbian weight $H=H_{neuron}+H_{astro}$.
@@
Astrocytes have a significant impact on the Hebbian weight due to their inherent non-linearity. A recent study by [17] introduced an AstroNet model that optimizes neural network connections through temporal and global regulation mechanisms. The authors found that astrocytes introduce non-linear effects, which compound the pre-existing nonlinearity of neurons. The concept of Hebbian Plasticity is associated with temporal and global regulation, where the current weight is influenced by past neuron weights and astrocytic weights. To map the effect of non-linearity, we consider that the summation of neuron and astrocyte Hebbian weights undergoes non-linear activation, ultimately generating the Hebbian weight ($H$). As such, we propose the integration of a sigmoid ($\sigma$) nonlinearity into the learning process. Thus, the weight $H$ exhibits the collective Hebbian plasticity of a tripartite synapse, as elucidated by Equation 11, where $K$ and $V$ correspond to keys and values of all the $N$ tokens in a single sequence (in matrix form: $K\in R^{N\times m}$, $V\in R^{N\times d}$). Assuming initial $H_{neuron,0}$ and $H_{astro,0}$ from Equations 9 and 10 are both zero, we can define the Hebbian weight in its matrix form ($H$) as,
$H=\sigma(H_{neuron}+H_{astro})\qquad\qquad\qquad\;\;\>$ (11)
$H=\sigma(\frac{1}{m}(\phi(K)^{T}V+W_{astro}*V))\qquad\quad\;\;$
Next, we consider formulation of the astrocytic activity parameter, $W_{astro}$. We note that astrocytic $Ca^{2+}$ dynamics is a relatively slower temporal process than the streaming rate of the input tokens. Therefore, it is expected that the astrocytic activity should encode relative information among successive tokens. In this context, we derive inspiration from algorithmic formulations of relative positional information of tokens, popularly used in Vision transformers [43, 44], and map it to the activity parameter $W_{astro}$. The astrocytic response parameter is therefore defined to use the edge of the tokens to capture relative position differences between input elements [43]. The edge of the tokens considers the pairwise relationships between input elements in the sense that the input is modeled as a labeled, directed, fully-connected graph. The edges pertaining to relative positional information are learnt using the temporal regulation property of astrocytes through the activity parameter, $W_{astro}$. Consequently, $H_{astro}$ can be considered as a representation of acquired positional knowledge derived from the input tokens. $W_{astro}$ can be modeled using Equation 12 where, $a_{ij}$ is the edge between input tokens $x_{i}$ and $x_{j}$.
@@ III-D2 Presynaptic Plasticity
The plasticity resulting from the induced calcium response inside the astrocyte due to bidirectional communication between the astrocyte and presynaptic neurons (hidden layer) can be defined as presynaptic plasticity. As tokens are sequentially presented to the hidden layer during the write mode, keys: $\phi(k_{t})$ are generated which can be mapped to an increasing number of released neurotransmitters (implying increment in $J_{channel}$ as per the computational neuroscience model), thereby resulting in elevation of the calcium concentration within the astrocyte (evident from Equation 5). Thus, presynaptic plasticity parameter $g$, can be mapped to calcium concentration in astrocytes modulated by indirect feedback ($eSP$), and can be expressed as the aggregate of the inputs ($h_{t}=\phi(k_{t})$) received by the presynaptic neuron. Equation 13 defines the formulation of the presynaptic plasticity parameter ($g$). Assuming the initial calcium concentration to be zero ($g_{0}=0$), $g$ can be expressed as,
@@
where, the exponent $\alpha$ encodes the degree of non-linearity. From our computational modelling, we use a value of $\alpha=0.25$. The subsequent section elaborates on the read mode operation which involves the process of retrieving the information encoded during the write mode to implement self-attention mechanism incorporating our key contributions towards more bio-plausible formulations of Hebbian and presynaptic plasticity.
@@ III-E Realization of Self-Attention
Self attention mechanism is realized when the encoded information is accessed during the read mode to obtain the activations of the output layer neurons. The hidden layer neurons receive queries ($q_{t}=x_{t}W_{Q}$) from the input layer and access the information that was previously encoded within the astrocyte. This information is stored via both the Hebbian weight ($H$) and the presynaptic plasticity parameter ($g$) during write mode. However, during read mode, $H$ is already learnt and not modulated by the transmitted queries from the presynaptic neurons as it is directly connected to the postsynaptic neurons (output layer). As $g$ is the presynaptic plasticity parameter, it has no direct connection to the postsynaptic neuron. Therefore, the information encoded in $g$ has to be retrieved by the queries transmitted from the hidden layer during read mode by inducing a calcium response from the astrocytes. In order to read the calcium response ($c_{t}$) of the astrocyte, the queries ($h_{t}$) are scaled by the corresponding presynaptic plasticity parameters ($g_{t}$) across all the $m$ hidden neurons and then summed together. This is depicted in Equation 15, where the matrix form of the calcium response is defined by $C$.
@@
Let us define the weight modulating the connection between the astrocyte and the output layer as $P$. According to studies regarding Tumor Necrosis Factor - $\alpha$ ($TNF-\alpha$), astrocytes possess the ability to boost synaptic weights in reaction to low neural activity and diminish weights in reaction to high neural activity [18] by means of calcium response (neural activity is proportional to released neurotransmitters). Consequently, it is possible to infer an inverse correlation between $P$ and the calcium response ($C$). Simplifying, this relation is given by,
$P=\frac{1}{\phi(Q)((\sum_{t=1}^{N}{\phi(k_{t})})^{\alpha})^{T}}$ (16)
@@
Given that the presynaptic plasticity modulated weight ($P$) is ascertained, we can define the network equations for the read mode. As seen from Fig. 2, the hidden layer and the output layer are connected by the Hebbian weight $H$ and the astrocyte is imposing another learnt weight (from calcium response) denoted by $P$ onto the existing weight $H$. Thus the weights are implemented by the Hadamard multiplication (element-wise: $\odot$) of $H$ and $P$ onto the output layer. Furthermore, the output layer receives the input ($F$) directly from the input layer modulated by unity weight to implement the residual connection present in the self-attention module of transformers. Considering the network architecture, the read mode network equations can be defined in Equation 17. It is important to highlight here that during the read mode, the decoding process occurs in a singular timestep, as opposed to the sequential nature of the write mode. Hence, all the parameters for input, hidden and output layers are represented in their matrix format.
$F=X\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\;\>$ (17)
$h=\phi(FW_{Q})=\phi(Q)\qquad\qquad\qquad\qquad\qquad\;$
$L=h\times(H\odot P)+F=\phi(Q)(H\odot P)+F\;\;\;$
@@
In Equation 17, all the parameters are in matrix mode, where the dimensions are defined as follows: $W_{Q}\in R^{d\times m}$, $X\in R^{N\times d}$ (input tokens) and $Q\in R^{N\times m}$ (queries). $L$ represents the generated output from the last layer. Self-attention algorithm is defined when the network completes a write and read mode as discussed next.
@@
By substituting the values of $H$ and $P$ from Equation 11 and 16 into Equation 17, the resulting Equation 18 can be obtained after the write and read mode:
$L=\frac{\phi(Q)(\sigma(\phi(K)^{T}V+W_{astro}V))}{m\times\phi(Q)\sum_{t=1}^{N}[{\phi(k_{t})}^{\alpha}]^{T}}+X$ (18)
In Equation 18, $L$ implements identical operation as the transformer self-attention from Equation 1 (along with the inclusion of an extra residual connection from the input layer). Thus, with this implementation, the astromorphic transformer is formulated.
@@
TABLE I: Performance comparison of our proposed astromorphic attention-based transformer architecture against other benchmarks with varying degrees of bioplausibility on the IMDB and CIFAR10 datasets.
Task
Model
Neuromorphic Implementation
Bioplausibility
Accuracy
@@
Sentiment Classification (IMDB Dataset)
Memristor SNN [45]
Directly trained in SNN domain
Yes
$84.86\%$
Memristor SNN [45]
Converted from ANN
Yes
$85.88\%$
ANN [45]
No spiking activity
No
$86.02\%$
Spike-transformer [46]
@@
Rate coded approach using Spiking Neurons
Yes
$86.36\%$
Recurrent SNN [47]
Presynaptic current based backpropagation
Yes
$86.82\%$
HRNaEMSA biGRU [48]
Implements multi-headed attention
No
$87.3\%$
HRNaEMSA biLSTM [48]
Implements multi-headed attention
No
$87.4\%$
CoRNN [49]
Coupled oscillators modeled as neurons
No
$87.4\%$
LSTM [50]
Memristor crossbars mimicking NN
No
$87.48\%$
SNN [51]
Compute-in-memory macro using SNN
Yes
$88.15\%$
UniCoRNN [52]
Solves vanishing gradient problem in BPTT
No
$88.4\%$
Transformer
Softmax attention without bioplausibilities
No
$88.9\pm 0.1\%$
Astromorphic Transformer (Our Model)
@@
Implements astromorphic attention with non-linearity and relative positional information
Yes
$\mathbf{88.7\pm 0.2\%}$
@@
Image Classification (CIFAR10 Dataset)
BioLeaF [53]
SNNs trained by BP-based rules
Yes
$86.88\%$
STBP in CIFARNET [54]
Spatio-temporal backprop
No
$89.83\%$
VGG16-SNN [55]
ANN-SNN conversion
Yes
$91.55\%$
MS-ResNet [56]
SNN oriented ResNet-110 architecture
No
$91.72\%$
ANN [57]
ResNet-19 (12.63M parameters)
No
$94.97\%$
ANN [57]
Transformer-4-384 (9.32M parameters)
No
$96.73\%$
Spikformer [57]
Spikformer-4-384 (9.32M parameters)
Yes
$95.19\%$
Spikingformer [54]
Spikingformer-4-384 (9.32M parameters)
Yes
$95.61\%$
Transformer
Softmax attention without bioplausibilities
No
$97.4\pm 0.1\%$
Astromorphic Transformer (Our Model)
@@
Implements astromorphic attention with non-linearity and relative positional information
Yes
$\mathbf{97.0\pm 0.2\%}$
Figure 4: Comparative analysis to examine the effects of incorporating non-linearity and relative positional encoding in the astromorphic transformer on the (a) IMDB and (b) CIFAR10 datasets. The yellow and purple lines in Fig. (b) are nearly overlapping.

### v3

@@ III-D1 Hebbian Plasticity
Hebbian plasticity is characterized by the sequential adjustment of synaptic weights in accordance with the interconnection pattern between two neurons (pre- and postsynaptic neurons) at the synapse. Previous studies [18] have exclusively focused on the plasticity occurring between two neurons, neglecting non-linearities and the plasticity that may occur between astrocytes and postsynaptic neurons, which is also involved in the learning process.
Contribution I: In the context of Hebbian plasticity, we underscore the importance of the addition of astrocyte-postsynaptic neuron connected Hebbian-inspired learning along with inherent non-linearities. To elaborate, the synaptic weights ($H_{neuron}$) that establish connections between the presynaptic and postsynaptic neurons are acquired through Hebbian plasticity during the write mode. This learning occurs due to the connection between pre- and postsynaptic neurons, which can be mapped to the synaptic current ($I_{syn}$) from the neuroscience perspective. As discussed previously, $I_{syn}$ is modulated by DSE, so $H_{neuron}$ is inspired from the direct feedback in the tripartite synapse. This leads to Hebbian plasticity between the hidden layer (presynaptic neuron) and the output layer (postsynaptic neuron). Hidden neuron outputs are denoted as $h_{t}=\phi(k_{t})$, whereas the output (last) layer outputs are denoted as $l_{t}=v_{t}$ in the write mode (refer to Equation 8). $H_{neuron}$ weights are learned via Equation 9, where $m$ acts as a scaling factor affecting the learning speed.
@@
The astrocyte functions as the third constituent in the tripartite synapse, enabling bidirectional communication with the synaptic neurons. Hence, we propose that the connection between the astrocyte and the postsynaptic neuron also be represented in this formulation through a learnable weight matrix ($H_{astro}$), in accordance with the Hebbian rule. Equation 10 shows the learning mechanism of these weights in which the presynaptic neuron output ($h_{t}$) is replaced by the astrocytic activity, represented by $w_{astro}$, as this connection incorporates the astrocyte and the postsynaptic neuron. The formulation of $w_{astro}$ will be considered in the succeeding text.
@@
The weight, denoted as $H_{astro}$, is inspired from the extrasynaptic current $SIC$ (refer to the tripartite synapse dynamics section). According to Equation 7, $SIC$ is added to $I_{syn}$ to generate the total synaptic current. Consequently, the two distinct learnable weights, denoted as $H_{neuron}$ and $H_{astro}$, are combined to produce the Hebbian weight $H=H_{neuron}+H_{astro}$.
@@
Astrocytes have a significant impact on the Hebbian weight due to their inherent non-linearity. A recent study by [17] introduced an AstroNet model that optimizes neural network connections through temporal and global regulation mechanisms. The authors found that astrocytes introduce non-linear effects, which compound the pre-existing non-linearity of neurons. The concept of Hebbian plasticity is associated with temporal and global regulation, where the current weight is influenced by past neuron weights and astrocytic weights. To map the effect of non-linearity, we consider that the summation of neuron and astrocyte Hebbian weights undergoes non-linear activation, ultimately generating the Hebbian weight ($H$). As such, we propose the integration of a sigmoid ($\sigma$) non-linearity into the learning process. Thus, the weight $H$ exhibits the collective Hebbian plasticity of a tripartite synapse, as elucidated by Equation 11, where $K$ and $V$ correspond to keys and values of all the $N$ tokens in a single sequence (in matrix form: $K\in R^{N\times m}$, $V\in R^{N\times d}$). Assuming initial $H_{neuron,0}$ and $H_{astro,0}$ from Equations 9 and 10 are both zero, we can define the Hebbian weight in its matrix form ($H\in R^{m\times d}$) as:
$H=\sigma(H_{neuron}+H_{astro})$ (11)
$H=\sigma(\frac{1}{m}(\phi(K)^{T}V+W_{astro}V))$
Next, we consider formulation of the astrocytic activity parameter $W_{astro}$ (Note, $w_{astro}$ in Equation 10 is the per-token representation and $W_{astro}$ in Equation 11 is the matrix representation). We note that astrocytic $Ca^{2+}$ dynamics is a relatively slower temporal process than the streaming rate of the input tokens. Therefore, it is expected that the astrocytic activity should encode relative information among successive tokens. In this context, we derive inspiration from algorithmic formulations of relative positional information of tokens, popularly used in Vision transformers [50, 51], and map it to the activity parameter $W_{astro}$. The astrocytic activity parameter is therefore defined to use the edge of the tokens to capture relative position differences between input elements [50]. The edge of the tokens considers the pairwise relationships between input elements in the sense that the input is modeled as a labeled, directed, fully connected graph. To model astrocytic activity, we first define the edges between consecutive input tokens $i$ and $j$ as $a_{ij}$ with the help of matrix $P\in R^{N\times N}$, containing the distance-based relative positional information between tokens, and $D\in R^{m\times m}$ which is a transformed version of $P$, achieved through a matrix basis transformation $D=MPM^{T}$. In this transformation, $M\in R^{m\times N}$ is a learnable matrix that projects the original $N$-dimensional space of positional relationships into an $m$-dimensional space. This transformation allows the model to learn an optimal representation of positional information, potentially reducing dimensionality if $m<N$, and capturing complex relationships between token positions. Mathematically, this transformation is a linear transformation, specifically a quadratic form, suitable for encoding pairwise interactions. Conceptually, this can be viewed as a learned change of basis, where $M$ defines a new coordinate system that optimizes the representation of positional information for the task. The resulting matrix $D$ encodes the spatial relative distances among tokens, represented as neuronal activity received by the astrocyte. Based on $D$, the edges are formulated as: $a_{ij}=W_{rel}D$, where $W_{rel}\in R^{N\times m}$ is a randomly initialized learnable weight matrix. To further enhance the efficiency of our model, we set $W_{rel}=M^{T}$. This design choice offers several advantages: First, it reduces the number of learnable parameters, promoting more efficient training and mitigating the risk of overfitting, especially in data-limited scenarios. Second, it encourages the learning of a more compact and generalizable representation, as the transformation $M$ now serves the dual purpose of both transforming the positional information ($P$ to $D$) and weighting the transformed information for edge computation ($a_{ij}=M^{T}D$). Finally, this constraint acts as a form of regularization, guiding the model to learn more robust and meaningful features by ensuring that the learned transformation is useful for both representing and processing positional relationships. The parameter, $m$, representing the number of neurons in the hidden layer, is crucial as it dictates the dimensionality of the learned representation. Its value is not fixed but rather can be adjusted based on the complexity of the dataset, allowing the model to adapt its capacity. Datasets with higher complexity may benefit from a larger $m$, providing the model with more representational power, while simpler datasets might learn effectively with a smaller $m$, promoting efficiency and potentially preventing overfitting. Building on the definition of edges, the astrocytic activity parameter $W_{astro}\in R^{m\times N}$ ($w_{astro}\in R^{m\times 1}$ for per token form) is defined as:
@@
where $\phi$ introduces a non-linearity to capture the temporal dynamics of astrocytes. While $W_{rel}$ is learned over time, the positional information provided by $D$ ensures that the model maintains awareness of relative distances among tokens, allowing $W_{astro}$ to adapt dynamically to the input. Consequently, $H_{astro}\in R^{m\times d}$ can be considered as a representation of acquired positional knowledge derived from the input tokens.
@@ III-D2 Presynaptic Plasticity
The plasticity resulting from the induced calcium response inside the astrocyte due to bidirectional communication between the astrocyte and presynaptic neurons (hidden layer) can be defined as presynaptic plasticity. As tokens are sequentially presented to the hidden layer during the write mode, keys: $\phi(k_{t})$ are generated, which can be mapped to an increasing number of released neurotransmitters (implying increment in $J_{channel}$ as per the computational neuroscience model), thereby resulting in elevation of the calcium concentration within the astrocyte (evident from Equation 5). Thus, the presynaptic plasticity parameter $g$ can be mapped to calcium concentration in astrocytes modulated by indirect feedback ($eSP$), and can be expressed as the aggregate of the inputs ($h_{t}=\phi(k_{t})$) received by the presynaptic neuron. Equation 13 defines the formulation of the presynaptic plasticity parameter ($g_{t},g\in R^{1\times m}$). Assuming the initial calcium concentration to be zero ($g_{0}=0$), $g$ can be expressed as:
@@
where the exponent $\alpha$ encodes the degree of non-linearity. From our computational modeling, we use a value of $\alpha=0.25$. The subsequent section elaborates on the read mode operation, which involves the process of retrieving the information encoded during the write mode to implement the self-attention mechanism incorporating our key contributions towards more bioplausible formulations of Hebbian and presynaptic plasticity.
@@ III-E Realization of Self-Attention
The self-attention mechanism is realized when the encoded information is accessed during the read mode to obtain the activations of the output layer neurons. The hidden layer neurons receive queries ($q_{t}=x_{t}W_{Q}$) from the input layer and access the information that was previously encoded within the astrocyte. This information is stored via both the Hebbian weight ($H$) and the presynaptic plasticity parameter ($g$) during write mode. However, during read mode, $H$ is already learned and not modulated by the transmitted queries from the presynaptic neurons, as it is directly connected to the postsynaptic neurons (output layer). As $g$ is the presynaptic plasticity parameter, it has no direct connection to the postsynaptic neuron. Therefore, the information encoded in $g$ has to be retrieved by the queries transmitted from the hidden layer during read mode by inducing a calcium response from the astrocytes. In order to read the calcium response ($c_{t}\in R^{1\times 1}$) of the astrocyte, the queries ($h_{t}$) are scaled by the corresponding presynaptic plasticity parameters ($g_{t}$) across all the $m$ hidden neurons and then summed together. This is depicted in Equation 15, where the matrix form of the calcium response is defined by $C\in R^{N\times 1}$.
@@
Let us define the weight modulating the connection between the astrocyte and the output layer as $P$. According to studies regarding Tumor Necrosis Factor - $\alpha$ ($TNF-\alpha$), astrocytes possess the ability to boost synaptic weights in reaction to low neural activity and diminish weights in reaction to high neural activity [18] by means of calcium response (neural activity is proportional to released neurotransmitters). Consequently, it is possible to infer an inverse correlation between $P$ and the calcium response ($C$). Simplifying, this relation ($P=\frac{1}{C}$) is given by Equation 16. To avoid the computational cost of a full matrix inversion, we compute the reciprocal calcium response element-wise by inverting only the non-zero elements of the diagonal matrix using a mask. This ensures numerical stability and computational efficiency for large-scale sequences.
$P=\frac{1}{\phi(Q)[(\sum_{t=1}^{N}{\phi(k_{t})})^{\alpha}]^{T}}$ (16)
@@
Given that the presynaptic plasticity modulated weight ($P$) is ascertained, we can define the network equations for the read mode. As seen from Fig. 2, the hidden layer and the output layer are connected by the Hebbian weight $H$, and the astrocyte is imposing another learned weight (from calcium response) denoted by $P$ onto the existing weight $H$. Thus the weights are implemented by the Hadamard (element-wise: $\odot$) multiplication of $H$ and $P$ onto the output layer. Furthermore, the output layer receives the input ($F$) directly from the input layer modulated by unity weight to implement the residual connection present in the self-attention module of transformers. Considering the network architecture, the read mode network equations can be defined in Equation 17. It is important to highlight here that during the read mode, the decoding process occurs in a singular timestep, as opposed to the sequential nature of the write mode. Hence, all the parameters for input, hidden, and output layers are represented in their matrix format (e.g., $X\in R^{N\times d}$, $h\in R^{N\times m}$, $L\in R^{N\times d}$ ).
$F=X$ (17)
$h=\phi(FW_{Q})=\phi(Q)$
$L=h(H\odot P)+F=\phi(Q)(H\odot P)+X$
@@
In Equation 17, all the parameters are in matrix mode, where the dimensions are defined as follows: $W_{Q}\in R^{d\times m}$ and $Q\in R^{N\times m}$ (queries). $L$ represents the generated output from the last layer. The self-attention algorithm is defined when the network completes a write and read mode, as discussed next.
@@
By substituting the values of $H$ and $P$ from Equations 11 and 16 into Equation 17, the resulting Equation 18 can be obtained after the write and read mode:
$L=\frac{\phi(Q)(\sigma(\phi(K)^{T}V+W_{astro}V))}{m\times\phi(Q)[(\sum_{t=1}^{N}{\phi(k_{t})})^{\alpha}]^{T}}+X$ (18)
$Y=\mathrm{Layernorm}(L)$
$\mathrm{logits}=\mathrm{Softmax}(\mathrm{Linear}(\mathrm{FFN}(Y)+Y))$
In Equation 18, $L$ implements an identical operation as the transformer self-attention from Equation 1 (along with the inclusion of an extra residual connection from the input layer). $Y$ is the layer-normalized version of $L$ and is further feedforwarded through fully connected linear layers to finally obtain the output probabilities in the form of logits. Thus, with this implementation, the Astromorphic Transformer is formulated. The detailed implementation of this process, including neuron-astrocyte interactions and synaptic modulation, is provided in Appendix E in the supplementary material, where the pseudo-code outlines the sequence of operations governing synaptic plasticity and calcium-mediated attention adjustments in the Astromorphic Transformer.
@@
TABLE I: Performance comparison of our proposed astromorphic attention-based transformer architecture against other benchmarks with varying degrees of bioplausibility on the IMDB, CIFAR10 and Wikitext-2 datasets.
Task Model Implementation Bio-plausibility Accuracy(%)/ Perplexity
@@
Sentiment Classification (IMDB Dataset) Memristor SNN [52] Directly trained in SNN domain Yes $84.86\%$
Memristor SNN [52] Converted from ANN Yes $85.88\%$
ANN [52] No spiking activity No $86.02\%$
@@
Spike-transformer [53] Rate-coded approach using spiking neurons Yes $86.36\%$
Recurrent SNN [54] Presynaptic current-based backpropagation Yes $86.82\%$
HRNaEMSA biGRU [55] Implements multi-headed attention No $87.3\%$
HRNaEMSA biLSTM [55] Implements multi-headed attention No $87.4\%$
CoRNN [56] Coupled oscillators modeled as neurons No $87.4\%$
LSTM [57] Memristor crossbars mimicking NN No $87.48\%$
SNN [58] Compute-in-memory macro using SNN Yes $88.15\%$
UniCoRNN [59] Solves vanishing gradient problem in BPTT No $88.38\%$
Vanilla Transformer Softmax attention without bioplausibilities No $88.9\pm 0.1\%$
Linearized Transformer inspired from astrocytes (Iso-architecture) [18] Linear attention with no astrocytic non-linearity or relative positional encoding Yes $88.4\pm 0.1\%$
@@
Astromorphic Transformer (Our Model) Implements astromorphic attention with non-linearity and relative positional information Yes $\mathbf{88.7\pm 0.2\%}$
@@
Image Classification (CIFAR10 Dataset) BioLeaF [60] SNNs trained by BP-based rules Yes $86.88\%$
STBP in CIFARNET [61] Spatio-temporal backprop No $89.83\%$
VGG16-SNN [62] ANN-SNN conversion Yes $91.55\%$
MS-ResNet [63] SNN oriented ResNet-110 architecture No $91.72\%$
ANN [64] ResNet-19 (12.63M parameters) No $94.97\%$
ANN [64] Transformer-4-384 (9.32M parameters) No $96.73\%$
Spikformer [64] Spikformer-4-384 (9.32M parameters) Yes $95.19\%$
Spikingformer [61] Spikingformer-4-384 (9.32M parameters) Yes $95.61\%$
Vanilla Transformer Softmax attention without bioplausibilities No $97.4\pm 0.1\%$
Linearized Transformer inspired from astrocytes (Iso-architecture)[18] Linear attention with no astrocytic non-linearity or relative positional encoding Yes $96.8\pm 0.2\%$
@@
Astromorphic Transformer (Our Model) Implements astromorphic attention with non-linearity and relative positional information Yes $\mathbf{97.0\pm 0.2\%}$
Language Modeling (Wikitext-2 Dataset) LSTM [67] Traditional RNN-based language model No $65.9$
AWD-LSTM [70] Standard LSTM with dropout, regularization No $58.8$
AWD-LSTM-MoS [71] Improved LSTM for language modeling No $63.88$
AWD-FWM [72] Fast weight memory-based language model No $61.65$
Vanilla Transformer Softmax attention without bioplausibilities No $73.4$
Linearized Transformer inspired from astrocytes (Iso-architecture) [18] Linear attention with no astrocytic non-linearity or relative positional encoding (does not converge due to gradient explosion) Yes $123.3$
Astromorphic Transformer (Our Model) Implements astromorphic attention with non-linearity and relative positional information Yes $\mathbf{33.8}$